I am having trouble with part B of this question.
let f(x)=sqrt(2x+6)
A. give the domain of f(x)
2x+6=> 0
2x=> -6
x=>-3
B. find F(a+1)I have no idea where to start with this part B? Any help would be appreciated. Thank you

Answer by AnlytcPhil(1810)

I am having trouble with part B of this question.
            ____ 
let f(x) = Ö2x+6
   A. give the domain of f(x)
2x+6=> 0
2x=> -6
x=>-3

B. find f(a+1)I have no idea where to start with this 
part B? Any help would be appreciated. Thank you
        ______
f(x) = Ö2x + 6

All you have to do is to replace all the x's by a + 1 
and simplify the right side:
            ______
    f(x) = Ö2x + 6
            ____________
f(a + 1) = Ö2(a + 1) + 6
            __________
f(a + 1) = Ö2a + 2 + 6
            ______
f(a + 1) = Ö2a + 8

Edwin